# Schrödinger symmetry: a historical review

### v1

@@ 1 Introduction
Symmetry [1] is a central concept in almost all theories of physical systems and their myriad applications. Symmetries arise either as internal symmetries or else as dynamical symmetries of time and space. Here, we are interested in the second class and notably in conformal invariance. Whenever it occurs, conformal invariance plays a crucial rôle in various theories and application, both for its mathematical and physical aspects. A necessary condition for conformal invariance is scale-invariance, and this requirement sharply distinguishes scale-invariant systems from those which do not have this property. Remarkably, scale-invariance does hold in certain conditions which happen to be of importance. Scale-invariance coupled to the usual symmetries of time and space, the conservation of energy-momentum and unitarity ‘normally’ leads to the emergence of the conformal group. [^8] For notable exceptions, see e.g. [18, 19, 20]. A systematic discussion on when scale-invariance does imply conformal invariance is given in [21, 22]. In high-energy physics, scale-invariance is significant in certain situations, for example in deep-inelastic scattering [2, 3] or more generally at the points of symmetry breaking, when it becomes exact and (for sufficiently local theories) gives rise to conformal field-theory (CFT) [4]. Conformal field-theory is among the main ingredients of string theory, see [5, 6] and refs. therein. On the low-energy side, conformal symmetry is useful in the description of equilibrium critical phenomena [7], most notably in two spatial dimensions [8], and the behaviour of physical systems near criticality [9].
Another development is to the AdS/CFT correspondence [10, 11] which is a conjectured relationship between two kinds of physical theories : it proposes a duality between theories in Anti-de Sitter space (AdS) and conformal field-theories on the boundary of AdS. This correspondence has been a subject of intense study in theoretical physics, in particular in string theory and quantum gravity.
@@
$\begin{array}[]{llcl}\hat{H}&=&\frac{m}{2}\dot{{\bm{x}}}^{2}&\text{\small Hamiltonian}\\[6.0pt] \hat{D}&=&\frac{1}{2}\big{(}{\bm{p}}\cdot{\bm{x}}+{\bm{x}}\cdot{\bm{p}}\big{)}-2t\hat{H}&\text{\small dilatation}\\[6.0pt] \hat{K}&=&\frac{m}{2}{{\bm{x}}}^{2}-tD-t^{2}\hat{H}&\text{\small expansions}\end{array}$ (1.3)
@@ 2 Carl Gustav Jacobi
$\bullet$ Motion of the centre of mass (CoM). Jacobi, in his Lecture 3, assumes that the “force function” only depends on the relative positions, $U=U({\bm{x}}^{a}-{\bm{x}}^{b})$, and derives what he calls “Das Princip131313‘Prinzip’ in present german orthography. der Erhaltung der Bewegung des Schwerpunkts” [Principle of the conservation of the motion of the centre of mass]. To this end, he chooses virtual displacements which correspond to shifting all positions by the same amount,
@@
providing us with an $N$-particle generalisation of the conserved quantity161616140 years later Niederer [44] derived dynamical symmetries for a particle with a time-dependent potential $V(t,{\bm{x}})$. In particular, if $V(t,{\bm{x}})=g(t)|{\bm{x}}|^{k}$, he found the presence of both dilatation and expansions for the case $k=-2$ and $g(t)$ constant [44, eq. (3.6)]. For the Newtonian potential with $k=-1$ the dilatation symmetry requires a time-dependent gravitational constant $g(t)$. But if the newtonian potential varies as $g(t)\sim t^{-1}$, expansions are a dynamical symmetry [44, eq. (3.8)]. On the other hand, if the potential has a time-dependence $g(t)\sim t^{-1/2}$, there is a dilatation dynamical symmetry [44, eq. (3.9)]. The first of those cases was observed independently in [75], consistently with an earlier suggestion of Dirac [76]. (2.5b) and we also recall the definition $\bm{p}_{a}=m_{a}\dot{{\bm{x}}}^{a}$ from (2.12).
@@ 3 Sophus Lie
The differential equation (3.33a) has two distinct characteristics and is then called hyperbolic. Lie shows again from his transformation theory that there is a canonical form $z_{xy}+Q(x-y)z_{y}+Z(x-y)z=0$ [27, eq. (10)]. Analysing the consequences of the infinitesimal point transformation
@@
where the functions $\xi,\eta,f$ are to be determined (and it is shown that $\varphi=0$) such that (3.33a) is transformed onto itself. The calculation is straightforward, if somewhat lengthy. It turns out that the most general solution reads [27, eq. (20)]
@@
where the functions $\xi,\eta,f$ are to be determined such that (3.33b) is transformed onto itself (and it is shown once more that $\varphi=0$). Straightforward calculations lead to a system of differential equations [27, eq. (27)] for $\xi,\eta,f,Z$. In particular, for $Z=0$ he finds
@@
such that the solution depends on the six parameters $\alpha,\beta,\gamma,m,n,\delta$ [^19] Notably, the parameters $\alpha,\beta,\gamma$ describe the ‘conformal’ transformations in the ‘time’ variable $y$, whereas $m,n$ describe spatial translations and Galilei-transformation in the ‘space’ variable $x$. Finally, $\delta$ describes a phase shift related to the central extension (called ‘mass’ in later sections). If $y$ is interpreted as time and $x$ as space, one recognises the infinitesimal transformations of the (centrally extended) Schrödinger group in $1+1$ time-space dimensions. Again, the solutions are written in form of the characteristics.
@@
It appears that both the sets of (ortho-)conformal transformations (3.35) as well as of Schröder-transformations (3.37) appear among the historically first examples of time-space dynamical symmetry transformations. In the remainder of Lie’s article [27], these symmetries are applied to reduce the solution of either the equations (3.33) to quadratures and a connection with the theory of minimal surfaces is pointed out. The possibility of an extension to more than two variables and/or to linear differential equations of higher order is mentioned. It must have looked to Lie too immediate to carry out explicitly, but was added by later generations of mathematicians [29, 30, 31].
@@
Table 1: Examples of infinite-dimensional groups of time-space transformations, defined through abstract coordinate transformations on time ($t$) and space coordinates ($\bm{r}\in\mathbb{R}^{d}$). A physical bias is parametrised by the constant $\beta\neq 0$ and distinguishes a preferred spatial direction $r_{\|}\in\mathbb{R}$ from transverse spatial directions $\bm{r}_{\perp}\in\mathbb{R}^{d-1}$. Time-space transformations are specified in terms of differentiable (vector-valued) functions $f,\bar{f},b,c,\bm{a}$ of their argument, $\dot{b}(t)={\rm d}b(t)/{\rm d}t$ and $\mathscr{R}\in\mbox{\sl SO}(d)$ is a rotation matrix. $\mathscr{S}$ is the invariant Schrödinger operator, where $\Delta_{\bm{r}}$ is the spatial laplacian.
This approach, based on (3.41), can be generalised to different Schrödinger operators. Table 1 lists examples of such symmetry transformations which can be extended, beyond the finite-dimensional sets considered by Jacobi, Lie and Niederer, to infinite-dimensional Lie algebras (which can be centrally extended). In particular, the infinitesimal conformal transformations and the Schrödinger-transformations already found by Lie and Jacobi are contained as the maximal finite-dimensional sub-algebras of the (ortho-)conformal and Schrödinger-Virasoro groups, along with the Schrödinger operator $\mathscr{S}$ on which they act as dynamical symmetries. Furthermore, given the explicit time-space transformations, it is easily checked that the full infinite-dimensional ortho- and meta-conformal transformation groups act as dynamical symmetries of their respective Schrödinger-operator, whereas for the Schrödinger-Virasoro and meta-Schrödinger-Virasoro groups, only the maximal finite-dimensional sub-group acts as dynamical symmetry group on $\mathscr{S}$.
@@ 5 Kastrup
A different approach to the dynamical symmetries of non-relativistic systems [91] deserves to be briefly mentioned. For a free non-relativistic particle, the kinetic energy $E=\frac{1}{2m}\bm{p}^{2}$. Hence its (non-constant) velocity is
@@
Defining $y^{0}:=vt$, with $v=|\bm{v}|$, one calls the set of points ${\tt y}=(y^{0},\bm{r})$ with the Minkowski metric ${\tt y}\cdot{\tt y}=\bigl{(}y^{0}\bigr{)}^{2}-\bm{r}^{2}$ the ‘Galilei space’. Kastrup explicitly gives the infinitesimal time-space transformations in Galilei space and states that their Lie algebra is isomorphic to the conformal algebra $\mathfrak{so}(2,4)$. The action integral $S=\int\!{\rm d}t\>\frac{1}{2m}\bigl{(}\frac{{\rm d}\bm{r}}{{\rm d}t}\bigr{)}^{2}$ is left invariant under these transformations. From this follow the conservation laws, including those from (1.3), of the non-relativistic free particle. This furnishes an example of a representation of the conformal algebra relevant for non-relativistic motion.
@@ 6 Jackiw-Niederer-Hagen-Barut
This extends the well-known Galilei symmetry by scale-invariance. Then the natural question arises whether this symmetry can further be extended to the conformal group by including also inversions [also called expansions]. We expect that conservation of energy-momentum plus scale-invariance require, just as in the relativistic theory, conformal invariance, – and this is indeed the case also in the non-relativistic limit [37]. An energy-momentum tensor may be explicitly constructed. It is traceless due to scale-invariance. Then an inversion operator can be constructed following ref. [37].
In conclusion, the free Schrödinger equation of a massive non-relativistic particle has, beyond the natural Galilean symmetry, two more “conformal” symmetries, cf. (1.3) or (2.5), respectively. Using Lie algebra language, the conserved quantities $D$ and $K$ in (1.3) close to an ${\rm o}(2,1)$ symmetry algebra with commutation relations,
@@
In “after-Noether” spirit, the operators (1.3) or their classical counterparts (2.5) are associated with the time-space transformations
@@
where $\tau,\lambda,\kappa$ are real parameters and, to distinguish from the Jacobi treatment of section 2, we introduced new, time-space coordinates $(T,{\bf X})$ as we shall discuss later.
@@
To understand the relation between Jacobi’s historic and the presently favoured time-space approaches we propose to view the transformations in (6.57) as a coordinate change. Restricting to just a single particle for simplicity ,
@@
Then Jacobi’s conserved quantity (2.27) becomes, with the notation (2.15) 262626${\mathcal{E}}={\mathcal{E}}^{CoM}$ for a one-particle theory., $D={D}\big{(}m{\bf X}^{2}/2\big{)}/{{\rm d}T}-2T{{\mathcal{E}}}/{\lambda^{2}}\,.$ But the energy scales also when we switch to the new coordinates
@@
Returning to the algebraic structure, we recall that the suggestion that the correct description a non-relativistic free quantum particle should involve the central extension of the Galilei group goes back to Inönü and Wigner [101] and to Bargmann [94]. In the ray representation of non-relativistic quantum mechanics, the plane-wave solution of the free Schrödinger equation transforms under the action of the Galilei group (in time and space)
@@
We finish this section with a comment. On p. 16 we recalled Niederer’s result that a Newtonian potential of the form
@@ 7 Local scale-invariance
As a consequence of much interest into equilibrium phase-transitions since the 1950s, conformal invariance was identified [7] as a key ingredient for the calculation of co-variant $n$-point correlation functions, [^29] Relabelling one of the spatial directions as ‘time’, a conformally covariant two-point function is $\bigl{\langle}\phi_{1}(t_{1},\bm{r}_{1})\phi_{2}(t_{2},\bm{r}_{2})\bigr{\rangle}=\delta_{x_{1},x_{2}}\bigl{(}(t_{1}-t_{2})^{2}+(\bm{r}_{1}-\bm{r}_{2})^{2}\bigr{)}^{-x_{1}}$, where $x_{1,2}$ are the scaling dimensions of the two scaling operators $\phi_{1,2}$. The only consequence of conformal invariance, beyond scale-invariance, is the constraint $x_{1}=x_{2}$. ‘Covariance’ means quasi-primary correlators [8]. especially in two spatial dimensions where the conformal algebra becomes infinite-dimensional and much stronger results hold than for $d>2$ [8]. One may wonder if Schrödinger transformations (or suitable extensions) can be considered similarly as generic time-space transformations such that $n$-point functions are fixed from the requirement of co-variance. To describe this, a more systematic notation is useful, see table 2. The Schrödinger algebra $\mathfrak{sch}(d)=\bigl{\langle}X_{\pm 1,0},{Y}_{\pm 1/2}^{(j)},M_{0},R^{(jk)}\bigr{\rangle}$ is written compactly as
@@
with $i,j,k,\ell=1,\ldots,d$. The generators in table 2 are for scalars under spatial rotations, see [37] for generalisations to higher spin. Eq. (7.73) immediately suggests an extension to the infinite-dimensional Schrödinger-Virasoro algebra [60] where $n,n^{\prime}\in\mathbb{Z}$, $m,m^{\prime}\in\mathbb{Z}+\frac{1}{2}$. An explicit representation of (7.73) through time-space transformation is (with $\bm{r}=(r_{1},\ldots,r_{d})$ and $\partial_{r_{j}}=\frac{\partial}{\partial r_{j}}$)
@@
Table 2: Different notations for the generators of the Schrödinger algebra in $d$ spatial dimensions ($j,k=1,\ldots,d$) and their definitions as time-space transformations.
Together with the conformal algebra, (7.73) forms the basis for the construction of more general time-space transformations with a prescribed dynamical scaling behaviour. [^30] From dynamical scaling with $z=2$ and time-/space-translation-invariance alone $\bigl{\langle}\phi_{1}(t_{1},\bm{r}_{1})\phi_{2}(t_{2},\bm{r}_{2})\bigr{\rangle}=(t_{1}-t_{2})^{-(x_{1}+x_{2})/2}\Phi\bigl{(}\frac{(\bm{r}_{1}-\bm{r}_{2})^{2}}{t_{1}-t_{2}}\bigr{)}$ with a non-trivial scaling function $\Phi$. One uses the following axioms [25] (for simplicity let $d=1$)
@@
These assumptions are sufficient to obtain a finite list of possibilities which can be stated as follows. The first result gives the form of $X_{n}$ as time-space transformations.
@@
$Y_{m}=Y_{k-N/2}=-\frac{2}{N(k+1)}\bigl{(}\frac{\partial a_{k}}{\partial r}\partial_{r}+\frac{\partial b_{k}}{\partial r}\bigr{)}$
@@
1. If $B_{1}=0$, then $\bigl{[}Y_{m},Y_{m^{\prime}}\bigr{]}=0$ and (7.79) is a closed Lie algebra, first identified by Negro, del Olmo and Rodríguez-Marco [82] and of which the special case $d=1$ is given in Theorem 2.343434Besides spatial rotations, in [82] there is a further purely spatial scaling generator $D_{s}$. It is nowadays referred to as $\ell$-conformal Galilei algebra. [^35] A lot of work has been dedicated to this class of algebras, focusing on possible central extensions and invariant equations [108, 109, 110, 111, 112] or physical realisations in the setting of the Pais-Uhlenbeck oscillator model [113, 114, 115, 116] or fluid mechanics [117]. However, for $B_{1}=0$ the generators do not contain the terms which through Ward identities would describe the transformation of scaling operators. Then one rests with time-space transformations which cannot be used to constrain $n$-point functions through their co-variance.The construction of such terms requires $B_{1}\neq 0$ but has not yet been solved satisfactorily, whenever $N\neq 1,2$. For generic $N$, it has been tried to use fractional (Riemann-Liouville) derivatives.
@@
3. Case (iii) gives for the generic situation $A_{1}\neq 0$ a Lie algebra isomorphism with the conformal algebra, but in a representation which does not conserve angles and is not conformal. While for a long time considered as a mere curiosity, it was only understood recently that this rather represents a new type of symmetry, the so-called meta-conformal algebra [83], which arises in systems with a directional bias in space. Their name alludes to the Lie algebra isomorphism with the standard ortho-conformal Lie algebra in $(1+1)$ time-space dimensions. In contrast to ortho-conformal invariance, in $(1+2)$ time-space dimensions there is an infinite-dimensional meta-conformal algebra, isomorphic to the direct sum of three Virasoro algebras. [^37] The same algebra also arises as dynamical symmetry of $1D$ spatially non-local erosion models [122, 123, 124, 125]. On the other hand, for $A_{1}=0$ one is back to the conformal galilean algebra. More will be said on this latter algebra in section 8.
Remarkably, these same infinite-dimensional Lie algebras can be recovered from the Newton-Cartan structures when an arbitrary anisotropy exponent $\theta$ is admitted [24]. These two symmetries arises when considering the geodesics of the Newton-Cartan structure. The end result is, in flat time-space
@@
This analysis centres on the time-space coordinate transformations and does not consider any Ward identities.
Figure 1: Spatial regions where various un-biased or biased symmetries can be realised. When distances scale isotropically with time as $r\sim\tau^{1/2}$, the time-space dynamical symmetry is the Schrödinger algebra. If a bias occurs and distances scale in the preferred direction as $r_{\|}\sim\tau$ while $r_{\perp}\sim\tau^{1/2}$ in the transverse direction, meta-Schrödinger invariance is realised. But if $r_{\|}\sim r_{\perp}\sim\tau$, meta-conformal symmetry may be realised, under certain conditions.

### v3

@@ 1 Introduction
Symmetry [1] is a central concept in almost all theories of physical systems and their myriad applications. Symmetries arise either as internal symmetries or else as dynamical symmetries of time and space. Here, we are interested in the second class and notably in conformal invariance. Whenever it occurs, conformal invariance plays a crucial rôle in various theories and application, both for its mathematical and physical aspects. A necessary condition for conformal invariance is scale-invariance, and this requirement sharply distinguishes scale-invariant systems from those which do not have this property. Remarkably, scale-invariance does hold in certain conditions which happen to be of importance. Scale-invariance coupled to the usual symmetries of time and space, the conservation of energy-momentum and unitarity ‘normally’ leads to the emergence of the conformal group. [^8] For notable exceptions, see e.g. [26, 27, 28]. A systematic discussion on when scale-invariance does imply conformal invariance is given in [29, 30]. In high-energy physics, scale-invariance is significant in certain situations, for example in deep-inelastic scattering [2, 3] or more generally at the points of symmetry breaking, when it becomes exact and (for sufficiently local theories) gives rise to conformal field-theory (CFT) [4]. Conformal field-theory is among the main ingredients of string theory, see [5, 6] and refs. therein. On the low-energy side, conformal symmetry is useful in the description of equilibrium critical phenomena [7], most notably in two spatial dimensions [8], and the behaviour of physical systems near criticality [9]. [^9] References on conformal invariance are legion. We limit ourselves to point to some well-known reviews/books which include [10, 11] and [12, 13, 14, 15, 16]. For a historical review see [17].
Another development deals with the AdS/CFT correspondence [18, 19] which is a conjectured relationship between two kinds of physical theories : it proposes a duality between theories in Anti-de Sitter space (AdS) and conformal field-theories on the boundary of AdS. This correspondence has been a subject of intense study in theoretical physics, in particular in string theory and quantum gravity.
@@
$\begin{array}[]{llcl}\hat{H}&=&\frac{m}{2}\dot{{\bm{x}}}^{2}~{}~{}=~{}~{}\frac{1}{2m}{\bm{p}}^{2}&\text{\small Hamiltonian}\\[6.0pt] \hat{D}&=&\frac{1}{2}\big{(}{\bm{p}}\cdot{\bm{x}}+{\bm{x}}\cdot{\bm{p}}\big{)}-2t\hat{H}&\text{\small dilatation}\\[6.0pt] \hat{K}&=&\frac{m}{2}{{\bm{x}}}^{2}-tD-t^{2}\hat{H}&\text{\small expansions}\end{array}$ (1.3)
@@ 2 Carl Gustav Jacobi
$\bullet$ Motion of the centre of mass (CoM). Jacobi, in his Lecture 3, assumes that the “force function” only depends on the relative positions, $U=U({\bm{x}}^{a}-{\bm{x}}^{b})$, and derives what he calls “Das Princip141414‘Prinzip’ in present German orthography. der Erhaltung der Bewegung des Schwerpunkts” [Principle of the conservation of the motion of the centre of mass]. To this end, he chooses virtual displacements which correspond to shifting all positions by the same amount,
@@
providing us with an $N$-particle generalisation of the conserved quantity171717140 years later Niederer [52] derived dynamical symmetries for a particle with a time-dependent potential $V(t,{\bm{x}})$. In particular, if $V(t,{\bm{x}})=g(t)|{\bm{x}}|^{k}$, he found the presence of both dilatation and expansions for the case $k=-2$ and $g(t)$ constant [52, eq. (3.6)]. For the Newtonian potential with $k=-1$ the dilatation symmetry requires a time-dependent gravitational constant $g(t)$. But if the Newtonian potential varies as $g(t)\sim t^{-1}$, expansions are a dynamical symmetry [52, eq. (3.8)]. On the other hand, if the potential has a time-dependence $g(t)\sim t^{-1/2}$, there is a dilatation dynamical symmetry [52, eq. (3.9)]. The first of those cases was observed independently in [83], consistently with an earlier suggestion of Dirac [84]. (2.5b) and we also recall the definition $\bm{p}_{a}=m_{a}\dot{{\bm{x}}}^{a}$ from (2.12).
@@ 3 Sophus Lie
The differential equation (3.33a) has two distinct characteristics and is then called hyperbolic. Lie shows again from his transformation theory that there is a canonical form $z_{xy}+Q(x-y)z_{y}+Z(x-y)z=0$ [35, eq. (10)]. Analysing the consequences of the infinitesimal point transformation
@@
where the functions $\xi,\eta,f$ are to be determined (and it is shown that $\varphi=0$) such that (3.33a) is transformed onto itself. The calculation is straightforward, if somewhat lengthy. It turns out that the most general solution reads [35, eq. (20)]
@@
where the functions $\xi,\eta,f$ are to be determined such that (3.33b) is transformed onto itself (and it is shown once more that $\varphi=0$). Straightforward calculations lead to a system of differential equations [35, eq. (27)] for $\xi,\eta,f,Z$. In particular, for $Z=0$ he finds
@@
such that the solution depends on the six parameters $\alpha,\beta,\gamma,m,n,\delta$ [^20] Notably, the parameters $\alpha,\beta,\gamma$ describe the ‘conformal’ transformations in the ‘time’ variable $y$, whereas $m,n$ describe spatial translations and Galilei-transformation in the ‘space’ variable $x$. Finally, $\delta$ describes a phase shift related to the central extension (called ‘mass’ in later sections). If $y$ is interpreted as time and $x$ as space, one recognises the infinitesimal transformations of the (centrally extended) Schrödinger group in $1+1$ space-time dimensions. Again, the solutions are written in form of the characteristics.
@@
It appears that both the sets of (ortho-)conformal transformations (3.35) as well as of Schrödinger-transformations (3.37) appear among the historically first examples of space-time dynamical symmetry transformations. In the remainder of Lie’s article [35], these symmetries are applied to reduce the solution of either the equations (3.33) to quadratures and a connection with the theory of minimal surfaces is pointed out. The possibility of an extension to more than two variables and/or to linear differential equations of higher order is mentioned. It must have looked to Lie too immediate to carry out explicitly, but was added by later generations of mathematicians [37, 38, 39].
@@
Table 1: Examples of infinite-dimensional groups of space-time transformations, defined through abstract coordinate transformations on time ($t$) and space coordinates ($\bm{r}\in\mathbb{R}^{d}$). A physical bias is parameterised by the constant $\beta\neq 0$ and distinguishes a preferred spatial direction $r_{\|}\in\mathbb{R}$ from transverse spatial directions $\bm{r}_{\perp}\in\mathbb{R}^{d-1}$. Time and space transformations are specified in terms of differentiable (vector-valued) functions $f,\bar{f},b,c,\bm{a}$ of their argument, $\dot{b}(t)={\rm d}b(t)/{\rm d}t$ and $\mathscr{R}\in\mbox{\sl SO}(d)$ is a rotation matrix. $\mathscr{S}$ is the invariant Schrödinger operator, where $\Delta_{\bm{r}}$ is the spatial Laplacian.
This approach, based on (3.41), can be generalised to different Schrödinger operators. Table 1 lists examples of such symmetry transformations which can be extended, beyond the finite-dimensional sets considered by Jacobi, Lie and Niederer, to infinite-dimensional Lie algebras (which can be centrally extended). In particular, the infinitesimal conformal transformations and the Schrödinger-transformations already found by Lie and Jacobi are contained as the maximal finite-dimensional sub-algebras of the (ortho-)conformal and Schrödinger-Virasoro groups, along with the Schrödinger operator $\mathscr{S}$ on which they act as dynamical symmetries. Furthermore, given the explicit space-time transformations, it is easily checked that the full infinite-dimensional ortho- and meta-conformal transformation groups act as dynamical symmetries of their respective Schrödinger-operator, whereas for the Schrödinger-Virasoro and meta-Schrödinger-Virasoro groups, only the maximal finite-dimensional sub-group acts as dynamical symmetry group on $\mathscr{S}$.
@@ 5 Kastrup
A different approach to the dynamical symmetries of non-relativistic systems [99, 17] deserves to be briefly mentioned. For a free non-relativistic particle, the kinetic energy $E=\frac{1}{2m}\bm{p}^{2}$. Hence its (non-constant) velocity is
@@
Defining $y^{0}:=vt$, with $v=|\bm{v}|$, one calls the set of points ${\tt y}=(y^{0},\bm{r})$ with the Minkowski metric ${\tt y}\cdot{\tt y}=\bigl{(}y^{0}\bigr{)}^{2}-\bm{r}^{2}$ the ‘Galilei space’. Kastrup explicitly gives the infinitesimal space-time transformations in Galilei space and states that their Lie algebra is isomorphic to the conformal algebra $\mathfrak{so}(2,4)$. The action integral $S=\int\!{\rm d}t\>\frac{m}{2}\bigl{(}\frac{{\rm d}\bm{r}}{{\rm d}t}\bigr{)}^{2}$ is left invariant under these transformations. From this follow the conservation laws, including those from (1.3), of the non-relativistic free particle. This furnishes an example of a representation of the conformal algebra relevant for non-relativistic motion.
@@ 6 Jackiw-Niederer-Hagen-Barut
This extends the well-known Galilei symmetry by scale-invariance. Then the natural question arises whether this symmetry can further be extended to the conformal group by including also expansions. We expect that conservation of energy-momentum plus scale-invariance require, just as in the relativistic theory, conformal invariance, – and this is indeed the case also in the non-relativistic limit [45]. An energy-momentum tensor may be explicitly constructed. It is traceless due to scale-invariance. Then an inversion operator can be constructed following ref. [45].
In conclusion, the free Schrödinger equation of a massive non-relativistic particle has, beyond the natural Galilean symmetry, two more “conformal” symmetries, cf. (1.3) or (2.5), respectively. Using Lie algebra language, the conserved quantities $D$ and $K$ in (1.3) close to an ${\rm so}(2,1)$ symmetry algebra with commutation relations,
@@
In “after-Noether” spirit, the operators (1.3) or their classical counterparts (2.5) are associated with the space-time transformations
@@
where $\tau,\lambda,\kappa$ are real parameters and, to distinguish from the Jacobi treatment of section 2, we introduced new, time and space coordinates $(T,{\bf X})$ as we shall discuss later.
@@
To understand the relation between Jacobi’s historic and the presently favoured space-time approaches we propose to view the transformations in (6.57) as a coordinate change. Restricting to just a single particle for simplicity ,
@@
Then Jacobi’s conserved quantity (2.27) becomes, with the notation (2.15) 272727${\mathcal{E}}={\mathcal{E}}^{CoM}$ for a one-particle theory., $D={D}\big{(}m{\bf X}^{2}/2\big{)}/{{\rm d}T}-2T{{\mathcal{E}}}/{\lambda^{2}}\,.$ But the energy scales also when we switch to the new coordinates
@@
Returning to the algebraic structure, we recall that the suggestion that the correct description of a non-relativistic free quantum particle should involve the central extension of the Galilei group goes back to Inönü and Wigner [109] and to Bargmann [102]. In the ray representation of non-relativistic quantum mechanics, the plane-wave solution of the free Schrödinger equation transforms under the action of the Galilei group (in time and space)
@@
We finish this section with a comment. On p. 17 we recalled Niederer’s result that a Newtonian potential of the form
@@ 7 Local scale-invariance
As a consequence of much interest into equilibrium phase-transitions since the 1950s, conformal invariance was identified [7] as a key ingredient for the calculation of co-variant $n$-point correlation functions, [^30] Relabelling one of the spatial directions as ‘time’, a conformally covariant two-point function is $\bigl{\langle}\phi_{1}(t_{1},\bm{r}_{1})\phi_{2}(t_{2},\bm{r}_{2})\bigr{\rangle}=\delta_{x_{1},x_{2}}\bigl{(}(t_{1}-t_{2})^{2}+(\bm{r}_{1}-\bm{r}_{2})^{2}\bigr{)}^{-x_{1}}$, where $x_{1,2}$ are the scaling dimensions of the two scaling operators $\phi_{1,2}$. The only consequence of conformal invariance, beyond scale-invariance, is the constraint $x_{1}=x_{2}$. ‘Covariance’ means quasi-primary correlators [8]. especially in two spatial dimensions where the conformal algebra becomes infinite-dimensional and much stronger results hold than for $d>2$ [8]. In this section, we give a brief and compact review on whether Schrödinger transformations (or suitable extensions) can be considered similarly as generic space-time transformations such that $n$-point functions are fixed from the requirement of co-variance. To describe this, a more systematic notation is useful, see table 2. The Schrödinger algebra $\mathfrak{sch}(d)=\bigl{\langle}X_{\pm 1,0},{Y}_{\pm 1/2}^{(j)},M_{0},R^{(jk)}\bigr{\rangle}$ is written compactly as
@@
with $i,j,k,\ell=1,\ldots,d$. The generators in table 2 are for scalars under spatial rotations, see [45] for generalisations to higher spin. As a first consequence, eq. (7.73) immediately suggests an extension to the infinite-dimensional Schrödinger-Virasoro algebra [68] where $n,n^{\prime}\in\mathbb{Z}$, $m,m^{\prime}\in\mathbb{Z}+\frac{1}{2}$. An explicit representation of (7.73) through space-time transformation is (with $\bm{r}=(r_{1},\ldots,r_{d})$ and $\partial_{r_{j}}=\frac{\partial}{\partial r_{j}}$)
@@
Table 2: Different notations for the generators of the Schrödinger algebra in $d$ spatial dimensions ($j,k=1,\ldots,d$) and their definitions as space-time transformations.
Together with the conformal algebra, (7.73) forms the basis for the construction of more general space-time transformations with a prescribed dynamical scaling behaviour. [^31] From dynamical scaling with $z=2$ and time-/space-translation-invariance alone $\bigl{\langle}\phi_{1}(t_{1},\bm{r}_{1})\phi_{2}(t_{2},\bm{r}_{2})\bigr{\rangle}=(t_{1}-t_{2})^{-(x_{1}+x_{2})/2}\Phi\bigl{(}\frac{(\bm{r}_{1}-\bm{r}_{2})^{2}}{t_{1}-t_{2}}\bigr{)}$ with a non-trivial scaling function $\Phi$. A simple way to specify this uses the following axioms [33] (for simplicity let $d=1$, but generalisations to $d>1$ appear straightforward)
@@
These assumptions are sufficient to obtain a finite list of possibilities which can be stated as follows. The first result gives the form of $X_{n}$ as space-time transformations.
@@
$Y_{m}=Y_{k-N/2}=-\frac{2}{N(k+1)}\left(\frac{\partial a_{k}}{\partial r}\partial_{r}+\frac{\partial b_{k}}{\partial r}\right)$
@@
Statements on complete algebras of dynamical symmetries can be given as follows, but only for the anisotropy exponent $\theta=1$ or $\theta=2$ can be Ward identities be specified directly from the symmetry generators.
@@
1. If $B_{1}=0$, then $\bigl{[}Y_{m},Y_{m^{\prime}}\bigr{]}=0$ and (7.79) is a closed Lie algebra, first identified by Negro, del Olmo and Rodríguez-Marco [90] and of which the special case $d=1$ is given in Theorem 2.353535Besides spatial rotations, in [90] there is a further purely spatial scaling generator $D_{s}$. It is nowadays referred to as $\ell$-conformal Galilei algebra. [^36] A lot of work has been dedicated to this class of algebras, focusing on possible central extensions and invariant equations [116, 117, 118, 119, 120] or physical realisations in the setting of the Pais-Uhlenbeck oscillator model [121, 122, 123, 124] or fluid mechanics [125]. However, for $B_{1}=0$ the generators do not contain the terms which through Ward identities would describe the transformation of scaling operators. Then one rests with space-time transformations which cannot be used to constrain $n$-point functions through their co-variance.The construction of such terms requires $B_{1}\neq 0$ but has not yet been solved satisfactorily, whenever $N\neq 1,2$. For generic $N$, it has been tried to use fractional (Riemann-Liouville) derivatives.
@@
3. Case (iii) gives for the generic situation $A_{1}\neq 0$ a Lie algebra isomorphism with the conformal algebra, but in a representation which does not conserve angles and is not conformal. While for a long time considered as a mere curiosity, it was only understood recently that this rather represents a new type of symmetry, the so-called meta-conformal algebra [91], which arises in systems with a directional bias in space. Their name alludes to the Lie algebra isomorphism with the standard ortho-conformal Lie algebra in $(1+1)$ space-time dimensions. In contrast to ortho-conformal invariance, in $(1+2)$ space-time dimensions there is an infinite-dimensional meta-conformal algebra, isomorphic to the direct sum of three Virasoro algebras. [^38] The same algebra also arises as dynamical symmetry of $1D$ spatially non-local erosion models [130, 131, 132, 133]. On the other hand, for $A_{1}=0$ one is back to the conformal Galilean algebra. More will be said on this latter algebra in section 8.
Remarkably, these same infinite-dimensional Lie algebras can be recovered from the Newton-Cartan structures when an arbitrary anisotropy exponent $\theta$ is admitted [32]. These two symmetries arises when considering the geodesics of the Newton-Cartan structure. The end result is, in flat space-time
@@
This analysis centres on the space-time coordinate transformations and does not consider any Ward identities.
Figure 1: Spatial regions where various un-biased or biased symmetries can be realised. When distances scale isotropically with time as $r\sim\tau^{1/2}$, the space-time dynamical symmetry is the Schrödinger algebra. If a bias occurs and distances scale in the preferred direction as $r_{\|}\sim\tau$ while $r_{\perp}\sim\tau^{1/2}$ in the transverse direction, meta-Schrödinger invariance is realised. But if $r_{\|}\sim r_{\perp}\sim\tau$, meta-conformal symmetry may be realised, under certain conditions.